OpenStudy (kaydra14):

Which sysyem has no solutions ? A. 2x + 5y = -3 -2x - 5y = -7 B. y = 3x 3x- y 0 C. 4x - y = 9 -4x - y = -1 D. 7x + 5y = 12 5x + 7y = 12

OpenStudy (anonymous):

Lines that do not intersect (parallel lines; having the same slope) have no solution. With that being said, try to graph each one and eliminate the answer choices that happen to intersect. (Desmos Graphing Calculator is very useful when it comes to solving systems of equations) :)

OpenStudy (kaydra14):

can you please see if i got the answer to this problem right please

OpenStudy (anonymous):

Of course! What did you get?

OpenStudy (kaydra14):

we are suppose to do linear combination and the equation was 5x - y = -21 x + y = -3 and the solution to te problem i got was ( -4 , -1 ) was i correct.

OpenStudy (anonymous):

Close! Watch your signs. The 1 should be positive. 5x -y = -21 x + y = -3 6x = -24 x = -4 x + y = -3 -4 + y = -3 +4 +4 y = 1 (-4, 1) Make sense? :)

OpenStudy (kaydra14):

Thanks now i understand. Can you check another one ?

OpenStudy (anonymous):

Np! Yep. :)

OpenStudy (kaydra14):

ok so the equations were 4x - 5y = 19 x + 5y = 36 and the solution to the problem was ( 11 , 5 ) . was i correct ?

OpenStudy (kaydra14):

@BunnyBree

OpenStudy (anonymous):

Yes! An easy way to check if you're correct is to plug your points back into the original equations. You could also graph them and see where they intersect.

OpenStudy (anonymous):

Did that help?

OpenStudy (kaydra14):

yes it did, but i am having trouble with one problem

OpenStudy (anonymous):

I can help. Wanna share?

OpenStudy (kaydra14):

3x + y = 4 2x + y = 5

OpenStudy (anonymous):

Okay. I'm gonna go ahead and do elimination. If I multiply one of the equations by -1, the y will become negative and I can easily eliminate y altogether. (3x + y = 4) (-1) 2x + y = 5 Now I have: -3x - y = -4 2x + y = 5 I'll solve for x. You can do y. :) -3x - y = -4 2x + y = 5 -x = 1 x = -1 Remember to plug x into one of the ORIGINAL equations.

OpenStudy (kaydra14):

so y = 7

OpenStudy (anonymous):

Yep. The solution is (-1, 7). :)